If an automobile gets an average of 27 miles per gallon on a trip and the standard deviation is 3 miles per gallon, find the probability that on a randomly selected trip, the automobile will get between 21 and 30 miles per gallon. Assume the variable is normally distributed.


1
Expert's answer
2021-03-24T14:21:43-0400

"P(21<X<30)"

"P(21<X<30)=P(\frac{21-27}{3}<Z<\frac{30-27}{3})"

"=P(-2<Z<1)"

"=P(Z<1)-P(Z<-2)"

From z tables, we have;

"0.84135-0.02286=0.8186"